OpenStudy (swag):

Distance = rate * time In the equation D = r * t, to solve for "r", one should: A.) Subtract "t" from both sides of the equation. B.) Divide both sides of the equation by "t". C.) Multiply both sides of the equation by "t". D.) Add a negative "t" to both sides of the equation.

Parth (parthkohli):

\[D = r\times t\]\[\implies {D \over t} = {r \times t \over t}\]Do you see what I am doing?

OpenStudy (anonymous):

i would set it up where the time cancles out leaving you with r=d/t

Parth (parthkohli):

Yes, so what did you exactly do there?